# CI/CD Configuration Practices in Open-Source Android Apps: An Empirical Study

### v1

@@ 4.3. RQ3: What are the common themes in CI/CD maintenance activities in Android apps?
Theme Validation via LDA. To further support the relevance and accuracy of themes identified in our manual analysis across all commits in our dataset, we perform LDA to generate topics automatically. The automatically generated topics by LDA align with our manually identified themes and provide broader insights into diverse CI/CD maintenance activities. Table 10 shows the $10$ topics generated by LDA, each with the most common keywords, an example commit message, and a brief description of each topic. We can see that the topics generated by LDA are consistent with the themes of the commit messages that we identify manually. For instance, topics #$\#1$ and #$\#5$ (with 25.5% of the commits) align with the Workflow/Pipeline Improvement and Build Fixing themes, despite the difference in proportion. Topics #$\#5$, #$\#6$, #$\#7$, and #$\#8$ (with 36.8% of the commits) correspond closely with the Environment Update, Tool Update, and Dependency Update themes, indicating a strong emphasis on keeping the build process up-to-date. We observe the reliance on the Renovate (9.2%) and Dependabot (6.9%) bots for automated dependency updates, while many other updates could potentially be performed manually. This makes it important for developers to be more aware of the availability and benefits of automated tools for various CI/CD improvements, and also signals to researchers the need to address the existing gaps in this context. With respect to testing and deployment, we observe topics #$\#2$, #$\#3$, #$\#9$, and #$\#10$ (with 37.7% of the commits) addressing these processes, with more focus toward improving testing (three topics with 31.5%) over deployment (only 6.2%). We further emphasize the critical role of deployment in mobile applications and recommend future research to explore the correlation between the frequency of deployment maintenance and the quality of delivered apps and user satisfaction, whether it leads to its enhancement or deterioration.
@@
Summary of RQ 3 Over a third of CI/CD configuration changes focus on improving workflow and fixing build issues, while another third target updates to the build environment, tools, and dependencies. The automatically generated topics align with our manually identified themes, offering broader insights into the diverse maintenance activities within CI/CD processes.
@@ 5. Implications
The need for standardized practices in CI/CD configurations. Our results in Observation 1.1 show a lack of standardized directives and values, indicating a need for uniform practices in configuration file conventions. This highlights the varied usage of CI/CD configuration practices, thus underlining the need for standardization. Though Java and Kotlin are found to share common directives across CI/CD services, this is not the case with Travis CI. This convergence in build configurations and deployment strategies promotes interoperability and standardization. The divergence in Travis CI reflects service-specific requirements or optimizations for these languages. This suggests that Java and Kotlin projects can leverage shared best practices, tools, or scripts. Therefore, for Travis CI, and also for other CI/CD services, developers may need to tailor configurations to specific service requirements, which also alarms the need for researchers to develop more automated approaches for that purpose.
Incorporating of CI/CD pipeline optimization mechanisms in practice. To support developers in adopting CI/CD practices, release management tools should offer automated solutions for configuring and optimizing CI/CD pipelines. In Observation 1.3, we find that projects use various app testing practices, including different testing tools and platforms, and the selection of appropriate emulators for their apps. However, traditional software testing and emulator use often pose delays to CI/CD pipelines. Despite extensive research on the selection and prioritization of devices for app testing (Khalid et al., 2014; Yang et al., 2024) or test cases (Elbaum et al., 2014; Pan et al., 2022), these approaches are not being applied in our studied projects. Successful adoption of advanced and recent testing mechanisms can help provide timely feedback to app developers about the quality (e.g., test coverage) of new app releases. Future research should focus more on developing automated testing frameworks that can easily align and integrate with existing CI/CD services and testing platforms to increase their chances of being adopted in practice. This may also include tools that automatically generate test cases or integrate test changes into the CI/CD pipeline.
@@
Recommendation mechanisms for updating CI/CD configurations. It is important to emphasize the critical role of automation tools, especially those that utilize bots, in maintaining up-to-date CI/CD pipelines. Our results in Observation 3.2 suggest an extensive use of bots, such as Dependabot, to automatically update software dependencies. It is also necessary to explore how these mechanisms can address broader CI/CD configuration updates, beyond mere dependency updates. Version control systems, such as GitHub, alongside CI/CD services should offer intelligent recommendation systems to autonomously propose configuration changes, aligning with the latest best practices and standards to ensures effective and efficient workflows. As Observation 3.1 indicates, developers often undertake numerous changes to improve the CI workflow and resolve build issues, which can span many stages of CI/CD pipelines. This gives an opportunity for advanced automated tools, leveraging generative AI and large language models, to systematically identify, implement, and refine improvements. Therefore, an investment in such intelligent tools is not only a pathway to minimizing manual interventions but also a strategy for keeping pace with innovative approaches in the ever-evolving realm of software development and integration. Researchers should also be part of this effort and should pay considerable attention to developing sophisticated techniques that align with the practical needs of software developers and build engineers.
More attention in testing and deployment phases of CI/CD pipelines. As Observation 2.4 demonstrates, while CI/CD pipeline maintenance activities heavily address setup and build stages, testing and deployment remain less frequently modified. This indicates a lack of continuous improvement in these areas. Despite their critical role in ensuring quality of mobile apps, testing and deployment phases appear relatively stable or unused, potentially due to manual intervention or insufficient automation tools. Future research should address this gap by exploring more advanced automated solutions, such as AI-driven test generation and optimized deployment strategies, by which developers can reduce manual overhead, improve feedback cycles, and enhance the overall effectiveness of their CI/CD pipelines. Furthermore, the industry could also benefit from these solutions if made open-source, as many organizations face challenges in investing in advanced automation tools due to budget constraints or lack of expertise.
@@ 6.1. Construct Validity
Construct threats to validity are concerned with the degree to which our analyses measure what we intend to analyze. In our study, we rely on data collected mostly from GitHub. Mistakenly computed values may affect our conclusions. To address this concern, we carefully filter and verify the data to minimize the possibility of computational errors that might impact our analyses. In addition, we use open coding to manually categorize the configuration directives into their respective phases or roles, which could be subjective. To mitigate this threat, multiple iterations of manual and automated analyses are performed by two co-authors of this paper to validate the generated categories using statistically significant samples of CI/CD configurations. Similarly, we perform a thematic analysis of commit messages to understand what maintenance activities developers undertake on CI/CD pipelines. Given that we are not the owners of the analyzed apps, there may be a misunderstanding of changes made in the studied commits. To address this, the analysis is also validated by two co-authors and further verified using automated topic modeling via LDA. Lastly, Finally, the reported results and statistics are calculated using our developed Python scripts. Though these scripts may contain defects that could lead to improper interpretations, we verify them by cross-checking the results with samples from our dataset.
@@ 6.2. Internal validity:
Internal threats to validity are concerned with the ability to draw conclusions from our empirical results. In our study, we investigate the association between project characteristics and maintenance activity in terms of CI/CD commit frequency using OLS linear regression. We recognize that our current variables may not fully capture all relevant factors, thus incorporating additional variables might change our conclusions and provide deeper insights, which we recommend for future research. In addition, we use LDA topic modeling for automatically identifying patterns in CI/CD commit messages. However, this could lead to generating incoherent or misleading topics due to poor parameter choices. To address these, we tune LDA parameters, such as the number of topics, alpha, and beta, applied data preprocessing, and manually refined the topic distributions. This ensures that our obtained results are meaningful, reliable, and aligned with our aim of validating manually identified themes of commit messages.
@@ 6.3. External validity:
External threats are concerned with our ability to generalize our study results. In our study, we initially collect 6,322 open-source Android apps that adopt CI/CD, relying on two sources that are commonly used by prior research: GitHub and F-Droid. When determining whether an app adopts CI/CD, we focus on well-known CI/CD services (e.g, GitHub Actions, Travis CI, CircleCI, and GitLab CI/CD) and their associated YML configuration files. We excluded CI/CD services for which we find only a limited number of adopting apps. However, some of the excluded apps may still use other CI/CD services that we did not account for. Also, some apps might not contain YML files, yet adopt CI/CD with default configurations or configured through web interfaces. Hence, our dataset might overlook some relevant apps, but those can be challenging to identify. Moreover, we realize that a considerable number of apps initially included in our dataset, which satisfied our filtration criteria, are not real apps but rather examples or samples provided as part of libraries or for educational purposes. We refer to these as toy projects. Therefore, to mitigate any possible threat to these apps, we perform an additional filtration step to exclude these types of apps from our dataset through rigorous analysis, resulting in a final set of 2,564 apps. Furthermore, our dataset is restricted to open-source Android apps, excluding non-free and iOS mobile apps, which limits the generalizability of our results. It is important to note that a major challenge in analyzing non-free apps is their limited public availability. However, our findings can still benefit developers using different CI/CD services, as they may encounter similar issues, such as dependency and testing issues. Future research should expand our study by exploring CI/CD adoption in non-free or iOS mobile apps and incorporating a wider range of CI/CD services.
@@ 7.1. Studies on analyzing mobile app releases
Differently from previous studies, this paper analyzes the challenges in CI/CD configurations that could potentially influence the above releasing issues. We explore how these configurations still rely on manual or external releasing mechanisms, which might contribute to deployment delays. We provide insights into optimizing release practices using automated techniques for better app store delivery.
@@ 7.2. Studies on CI/CD adoption and configurations
Martin Fowler proposed ten best practices for adopting CI in software projects (Fowler, [n.d.]), which researchers then studied to understand the impact of these practices in real-world projects (Zhao et al., 2017). Hilton et al. (2016b) analyzed thousands of GitHub projects and surveyed 442 software developers on CI adoption challenges, and discovered that only 40% of the projects use CI, with the primary barrier being the lack of experience of developers. Widder et al. (2019) also identified CI adoption challenges by reviewing 37 studies and surveying 132 developers, and reported common difficulties in troubleshooting, inconsistency of tools across CI services, and complex setups. Elazhary et al. (2021) also studied the challenges in CI adoption, and found that automating UI testing and pull request review bottlenecks represent the main issue. Zampetti et al. (2020) developed a catalog of bad CI/CD practices from developer interviews and Stack Overflow analysis, highlighting mismanaged branching and versioning, and unaddressed infrastructure, configuration, and testing smells. Vassallo et al. (2020) studied CD configuration issues and developed a tool to detect relevant configuration smells.
@@
Renggli et al. (2019) examined the unique challenges of applying CI to machine learning projects, highlighting the need for specialized tools and workflows to address issues absent in regular software development. Rzig et al. (2022) also investigated CI adoption issues, finding lower CI adoption rates in ML projects compared to other software. They identified that common CI challenges, like debugging build failures, are worsened by ML’s specific needs, such as large datasets.
@@ 7.3. Studies on CI/CD for mobile apps
Klepper et al. (2015) proposed the Rugby process as an agile process with support for continuous delivery. They applied the Rugby process to the different case studies of mobile apps. Rossi et al. (2016) demonstrate their experience of adopting the testing and continuous deployment practices in Facebook. Later, Cruz et al. (2019a) studied testing practices in open-source apps. They found that continuous integration and continuous deployment pipelines are not widely adopted in the studied mobile apps. Recently, Wang et al. (2023b) studied the adoption of UI testing in 318 Google Play Store apps deployed on GitHub. Though they observed that 40% of the studied apps use CI services (e.g., GitHub Actions and Travis CI), they found that the adoption of CI/CD services is primarily used for unit testing. Similar to our study, Liu et al. (2022) analyzed 7,899 Android apps hosted on GitHub, Bitbucket, and GitLab that adopt CI/CD services, with GitHub Actions and Travis CI as the most commonly adopted. However, their work mainly offers general statistics about the adoption of CI/CD services without delving deeper into specific configurations or maintenance practices across CI/CD services. Furthermore, their dataset was not rigorously filtered to ensure that only actual Android apps were included.
This paper complements previous research by conducting more in-depth analyses of CI/CD pipeline configuration practices and maintenance activities. In contrast to the above studies, our investigation span across four different CI/CD services, namely GitHub Actions, Travis CI, CircleCI, and GitLab CI/CD, on a carefully curated set of $2.5k$ Android apps collected from GitHub and F-Droid. This enabled us to uncover the absence of commonalities among these services and identify potential trade-offs when selecting a CI/CD service for a specific project.

### v2

@@ 4.3. RQ3: What are the common themes in CI/CD maintenance activities in Android apps?
Theme Validation via LDA. To further support the relevance and accuracy of themes identified in our manual analysis across all commits in our dataset, we perform LDA to generate topics automatically. The automatically generated topics by LDA align with our manually identified themes and provide broader insights into diverse CI/CD maintenance activities. Table 10 shows the $10$ topics generated by LDA, each with the most common keywords, an example commit message, and a brief description of each topic. We can see that the topics generated by LDA are consistent with the themes of the commit messages that we identify manually. For instance, topics #$\#1$ and #$\#5$ (with 25.5% of the commits) align with the Workflow/Pipeline Improvement and Build Fixing themes, despite the difference in proportion. Topics #$\#5$, #$\#6$, #$\#7$, and #$\#8$ (with 36.8% of the commits) correspond closely with the Environment Update, Tool Update, and Dependency Update themes, indicating a strong emphasis on keeping the build process up-to-date. We observe the reliance on the Renovate (9.2%) and Dependabot (6.9%) bots for automated dependency updates, while many other updates could potentially be performed manually. This makes it important for developers to be more aware of the availability and benefits of automated tools for various CI/CD improvements, and also signals to researchers the need to address the existing gaps in this context. With respect to testing and deployment, we observe topics #$\#2$, #$\#3$, #$\#9$, and #$\#10$ (with 37.7% of the commits) addressing these processes, with more focus toward improving testing (three topics with 31.5%) over deployment (only 6.2%). We further emphasize the critical role of deployment in mobile apps and recommend future research to explore the correlation between the frequency of deployment maintenance and the quality of delivered apps and user satisfaction, to determine whether it leads to enhancement or deterioration.
@@
Summary of RQ 3 Over a third of CI/CD configuration changes focus on improving workflow and fixing build issues, while another third target updates to the build environment, tools, and dependencies. The automatically generated topics align with our manually identified themes, offering broader insights into the diverse maintenance activities within CI/CD processes. This is consistent with some prior findings on general-purpose projects that highlight workflow restructuring and environment updates as common CI change motivations, but still inconsistent with other findings that report more frequent changes related to documentation, notifications, and deployment than we observed in Android projects.
@@ 5. Implications
The need for standardized practices in CI/CD configurations. Our results in Observation 1.1 show a lack of standardized directives and values, indicating a need for uniform practices in configuration file conventions. This highlights the varied usage of CI/CD configuration practices, thus underlining the need for standardization. Though Java and Kotlin are found to share common directives across CI/CD services, this is not the case with Travis CI reflects service-specific requirements or optimizations for these languages. This suggests that Java and Kotlin projects can leverage shared best practices, tools, or scripts. Therefore, for any CI/CD service, developers may need to tailor configurations to specific service requirements, which also alarms the need for researchers to develop more automated approaches for that purpose. Building on this, we argue that standardized practices—such as improved Android-specific documentation, curated configuration templates, reusable components like GitHub Actions workflow templates [^28] https://docs.github.com/en/actions/writing-workflows/using-workflow-templates and CircleCI orbs [^29] https://circleci.com/orbs, and Android-focused best-practice guidelines—can help address the lack of consistency observed across services. These practices can collectively contribute to reducing configuration effort, minimizing errors, improving maintainability, ensuring consistency, and simplifying migration across CI/CD tools, ultimately lowering the barrier to effective CI/CD adoption in Android development. Moreover, new developers or team members may face difficulties in understanding and contributing to projects with varying CI/CD configurations, leading to longer onboarding times and challenges in knowledge transfer. These factors can negatively affect team efficiency and productivity, further reinforcing the need for standardized CI/CD practices, especially in the context of Android app development.
Incorporating CI/CD pipeline optimization mechanisms in practice. To support developers in adopting CI/CD practices, release management tools should offer automated solutions for configuring and optimizing CI/CD pipelines. In Observation 1.3, we find that projects use various app testing practices, including different testing tools and platforms, and the selection of appropriate emulators for their apps. However, traditional software testing and emulator use often pose delays to CI/CD pipelines. Despite extensive research on the selection and prioritization of devices for app testing (Khalid et al., 2014; Yang et al., 2024) or test cases (Elbaum et al., 2014; Pan et al., 2022), these approaches are not being applied in our studied projects. Successful adoption of advanced and recent testing mechanisms can help provide timely feedback to app developers about the quality (e.g., test coverage) of new app releases. Future research should focus more on developing automated testing frameworks that can easily align and integrate with existing CI/CD services and testing platforms to increase their chances of being adopted in practice. This may also include tools that automatically generate test cases or integrate test changes into the CI/CD pipeline.
@@
Recommendation mechanisms for updating CI/CD configurations. It is important to emphasize the critical role of automation tools, especially those that utilize bots, in maintaining up-to-date CI/CD pipelines. Our results in Observation 3.2 suggest an extensive use of bots, such as Dependabot, to automatically update software dependencies. It is also necessary to explore how these mechanisms can address broader CI/CD configuration updates, beyond mere dependency updates. Version control systems, such as GitHub, alongside CI/CD services, should offer intelligent recommendation systems to autonomously propose configuration changes, aligning with the latest best practices and standards to ensure effective and efficient workflows. As Observation 3.1 indicates, developers often undertake numerous changes to improve the CI workflow and resolve build issues, which can span many stages of CI/CD pipelines. This gives an opportunity for advanced automated tools, leveraging generative AI and large language models, to systematically identify, implement, and refine improvements. Therefore, an investment in such intelligent tools is not only a pathway to minimizing manual interventions but also a strategy for keeping pace with innovative approaches in the ever-evolving realm of software development and integration. Researchers should also be part of this effort and should pay considerable attention to developing sophisticated techniques that align with the practical needs of software developers and build engineers.
More attention in testing and deployment phases of CI/CD pipelines. As Observation 2.4 demonstrates, while CI/CD pipeline maintenance activities primarily address setup and build processes, testing and deployment remain less frequently modified, indicating limited continuous improvement in these critical phases. This gap may be partly due to manual intervention or a lack of robust automation tools. The challenge is even greater in Android apps, where testing often depends on emulators to support a wide range of devices and OS versions. These requirements complicate integration with third-party services, such as Google Play Console [^30] https://play.google.com/consoleand AWS Device Farm [^31] https://aws.amazon.com/device-farm, creating further barriers to automation. Deployment adds another layer of complexity, as Android apps must be pushed to app stores like the Google Play Store, where updates undergo review processes involving security, compatibility, and privacy checks, unlike web applications, which can be updated directly. Given these constraints, CI/CD services such as GitHub Actions often recommend performing app deployment manually and externally from the main CI/CD workflow. Therefore, future research and CI/CD service providers should explore advanced automated solutions, such as AI-driven test generation (Wang et al., 2024; Su et al., 2024; Schäfer et al., 2023) and optimized deployment strategies (Vadde and Munagandla, 2022), to reduce manual overhead, improve feedback cycles, and enhance the overall effectiveness of mobile CI/CD pipelines. Promising directions include automated test case generation tailored to CI/CD pipelines beyond prompt engineering that relies solely on the latest error messages, but is rather based on log histories and previous fixes (Tufano et al., 2019). Example tools include Launchable [^32] https://www.launchableinc.com and Diffblue [^33] https://www.diffblue.com, neither of which is adopted in our dataset. In terms of deployment, to enhance existing support by platforms like GitHub in generating release notes [^34] https://docs.github.com/en/repositories/releasing-projects-on-github/automatically-generated-release-notes, AI could further help in release readiness assessment (Hoang et al., 2020) and detection of compliance risks by analyzing code changes and issue history (Yang et al., 2025). These capabilities could be integrated into CI/CD platforms as reusable workflows to allow adaptation to project-specific configurations over time. Despite the existence of paid AI-driven tools for software deployment, such as Digital.ai [^35] https://digital.ai, and research-based tools, such as LADs (Faraz Khan et al., 2025), our dataset does not show any utilization of such tools. By integrating AI in existing tools and research, these directions offer practical and impactful ways to reduce manual effort, enhance reliability, and standardize CI/CD practices across Android projects. Furthermore, open-sourcing such tools could further benefit the software engineering community, particularly for teams with limited budgets or expertise.
@@ 6.1. Construct Validity
Construct threats to validity are concerned with the degree to which our analyses measure what we intend to analyze. In our study, we rely on data collected mostly from GitHub and analyzed using Python scripts written by the authors of this paper. Mistakenly computed values may affect our conclusions. To address this concern, we carefully filter and verify the data to minimize the possibility of computational errors that might impact our analyses. In addition, we use open coding to manually categorize the configuration directives into their respective phases or roles, which could be subjective. To mitigate this threat, multiple iterations of manual and automated analyses are performed by two co-authors of this paper with practical experience in software engineering and CI/CD practices to validate the generated categories using statistically significant random samples of CI/CD configurations. This process of validation, refinement, and re-evaluation was repeated iteratively six times for the remaining unfiltered projects until we reached an observed agreement of $93\%$ in manually validating the automatically filtered projects, ensuring that our final dataset reliably excluded toy projects and retained real Android mobile apps. Similarly, we perform a thematic analysis of commit messages to understand what maintenance activities developers undertake on CI/CD pipelines. Given that we are not the owners of the analyzed apps, there may be a misunderstanding of changes made in the studied commits. To address this, the analysis is also validated by two co-authors and further verified using automated topic modeling via LDA. Lastly, the reported results and statistics are calculated using our developed Python scripts. Though these scripts may contain defects that could lead to improper interpretations, we verify them by cross-checking the results with samples from our dataset.
@@ 6.2. Internal validity:
Internal threats to validity are concerned with the ability to draw conclusions from our empirical results. In our study, we investigate the association between project characteristics and maintenance activity in terms of CI/CD commit frequency using OLS linear regression. We recognize that our current variables may not fully capture all relevant factors, thus incorporating additional variables might change our conclusions and provide deeper insights, which we recommend for future research. In addition, we use LDA topic modeling for automatically identifying patterns in CI/CD commit messages. However, this could lead to generating incoherent or misleading topics due to poor parameter choices. To address these, we perform data preprocessing, tune LDA parameters (such as the number of topics, alpha, and beta), and manually refine the topic distributions. This ensures that our obtained results are meaningful, reliable, and aligned with our aim of validating manually identified themes of commit messages.
@@ 6.3. External validity:
External threats are concerned with our ability to generalize our study results. In our study, we initially collect 6,322 open-source Android apps that adopt CI/CD, relying on two sources that are commonly used in prior research: GitHub and F-Droid. When determining whether an app adopts CI/CD, we focus on well-known CI/CD services (e.g, GitHub Actions, Travis CI, CircleCI, and GitLab CI/CD) and their associated YML configuration files. We excluded CI/CD services for which we find only a limited number of adopting apps. However, some of the excluded apps may still use other CI/CD services that we did not account for. Also, some apps might not contain YML files, yet adopt CI/CD with default configurations or configured through web interfaces. Hence, our dataset might overlook some relevant apps, but those can be challenging to identify. Moreover, we realize that a considerable number of apps initially included in our dataset, which satisfied our filtration criteria, are not real apps but rather examples or samples provided as part of libraries or for educational purposes. We refer to these as toy projects. Therefore, to mitigate any possible threat to these apps, we perform an additional filtration step to exclude these types of apps from our dataset through rigorous analysis, resulting in a final set of 2,557 apps. Furthermore, our dataset is restricted to open-source Android apps, excluding non-free, iOS, and non-Android apps, which limits the generalizability of our results. A major challenge in analyzing non-free apps is their limited public availability. In addition, the lack of comparisons with non-Android projects prevents us from drawing broader conclusions about CI/CD adoption, configuration complexity, and maintenance practices across domains. To partially address this threat, we compare and contrast our findings with prior research findings on general-purpose CI/CD practices, allowing us to situate our observations within broader trends. Nevertheless, our findings may still benefit developers using various CI/CD services, as common challenges, such as dependency management and testing, are widely encountered. Therefore, we aim to further expand our analyses and encourage future research to include a broader range of application domains to support more generalizable conclusions and to explore domain-specific patterns in CI/CD configuration and maintenance practices.
Our study focuses exclusively on Android apps, which may limit the generalizability of our conclusions across other application domains. While this focus allows for a deep investigation of mobile-specific CI/CD practices, we acknowledge that comparisons with non-Android projects, particularly regarding CI/CD adoption, configuration complexity, and maintenance frequency, could strengthen our findings. We aim to expand our analyses in future work to include a broader range of application types to support more generalizable conclusions.
@@ 7.1. Studies on analyzing mobile app releases
Unlike previous studies, this paper analyzes the challenges in CI/CD configurations that could potentially influence the above releasing issues. We explore how these configurations still rely on manual or external releasing mechanisms, which might contribute to deployment delays. We provide insights into optimizing release practices using automated techniques for better app store delivery.
@@ 7.2. Studies on CI/CD adoption and configurations
Martin Fowler proposed ten best practices for adopting CI in software projects (Fowler, [n.d.]), which researchers then studied to understand the impact of these practices in real-world projects (Zhao et al., 2017). Hilton et al. (2016b) analyzed thousands of GitHub projects and surveyed 442 software developers on CI adoption challenges, and discovered that only 40% of the projects use CI, with the primary barrier being the lack of experience of developers. Widder et al. (2019) also identified CI adoption challenges by reviewing 37 studies and surveying 132 developers, and reported common difficulties in troubleshooting, inconsistency of tools across CI/CD services, and complex setups. Elazhary et al. (2021) also studied the challenges in CI adoption, and found that automating UI testing and pull request review bottlenecks represent the main issue. Zampetti et al. (2020) developed a catalogue of bad CI/CD practices from developer interviews and Stack Overflow analysis, highlighting mismanaged branching and versioning, and unaddressed infrastructure, configuration, and testing smells. Vassallo et al. (2020) studied CD configuration issues and developed a tool to detect relevant configuration smells.
@@
Rzig et al. (2022) also investigated CI adoption issues, finding lower CI adoption rates in ML projects compared to other software. They identified that common CI challenges, like debugging build failures, are worsened by ML’s specific needs, such as large datasets.
@@ 7.3. Studies on CI/CD for mobile apps
Klepper et al. (2015) proposed the Rugby process as an agile process with support for continuous delivery. They applied the Rugby process to the different case studies of mobile apps. Rossi et al. (2016) demonstrate their experience of adopting the testing and continuous deployment practices in Facebook. Later, Cruz et al. (2019a) studied testing practices in open-source apps. They found that continuous integration and continuous deployment pipelines are not widely adopted in the studied mobile apps. Recently, Wang et al. (2023b) studied the adoption of UI testing in 318 Google Play Store apps deployed on GitHub. Though they observed that 40% of the studied apps use CI/CD services (e.g., GitHub Actions and Travis CI), they found that the adoption of CI/CD services is primarily used for unit testing. Similar to our study, Liu et al. (2022) analyzed 7,899 Android apps hosted on GitHub, Bitbucket, and GitLab that adopt CI/CD services, with GitHub Actions and Travis CI as the most commonly adopted. However, their work mainly offers general statistics about the adoption of CI/CD services without delving deeper into specific configurations or maintenance practices across CI/CD services. Furthermore, their dataset was not rigorously filtered to ensure that only actual Android apps were included.
This paper complements previous research by conducting more in-depth analyses of CI/CD pipeline configuration practices and maintenance activities. In contrast to the above studies, our investigation spans across four different CI/CD services, namely GitHub Actions, Travis CI, CircleCI, and GitLab CI/CD, on a carefully curated set of $2.5k$ Android apps collected from GitHub and F-Droid. This enabled us to uncover the absence of commonalities among these services and identify potential trade-offs when selecting a CI/CD service for a specific project.